Operations Report – April 18th

Crew 192 Operations Report 18Apr2018

SOL 10

Name of person filing report: Victoria Varone

Non-nominal systems: None

Generator (hours run): 12 hours

Solar ‒ SOC 66%

Diesel – 75% fuel

Propane – 62% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – All ATVs filled 100%, 25% left in one can, requesting gas refill.

Water (trailer) – Empty, tank being cleaned in RAM.
Water (static) – 690 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133669.6

Toilet tank emptied: No. Last emptied April 16.

ATV’s Used: None

Oil Added? No but checked and OK.

ATV Fuel Used: 0% today

# Hours the ATVs were Used today: 0

Notes on ATVs: ATVs fine, oil checked, gas filled yesterday.

Deimos rover used: Yes (out of commission per Shannon)
Hours: 121.5
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: Yes
Hours: 32.8 (corrected)
Beginning charge: 100%
Ending charge: 68%
Currently charging: Yes

Opportunity rover used: yes
Hours: 29.8
Beginning charge: 100%
Ending charge: 81%
Currently charging: Yes

Curiosity rover used: Yes
Hours: 27.4
Beginning charge: 100%
Ending charge: 67%
Currently charging: No

HabCar used and why, where? No

General notes and comments: Some great EVAs on a beautiful day.

Summary of internet: 0MB at 1900.

Summary of suits and radios: Suits all charging, three one-piece suits used, radios working fine.

Summary of Hab operations: Cooked, cleaned, and organized the hab.

Summary of GreenHab operations: See GreenHab Report.

Summary of ScienceDome operations: None

Summary of RAM operations: Trailer water tank remains in the RAM to air out after being cleaned for the final time yesterday, will be checked tomorrow.

Summary of health and safety issues: None

Questions, concerns, and requests to Mission Support: – Hab upper deck faucet hot water shut due to increasing leaks.

– Fridge got hot last night despite being closed and plugged in properly, was unplugged and plugged back and is working fine again.

– Steps in front of rear airlock were secured again after become a bit loose after the first fix.

Sol Summary – April 18th

Crew 192 Sol Summary Report 18Apr2018

Sol 10

Summary Title: Canyons, Baby

Author: Victoria Varone

Mission Status: A beautiful day for the crew, mission nearing its end.

Sol Activity Summary: The crew went on two amazing EVAs today, exploring some of the major Martian canyons in the area. Lith Canyon and Murphy Canyon provided some of the most breathtaking views we’ve seen so far.

Then we get a little sun-gazing in at the Musk Observatory before cooking an awesome dinner for the crew that was a wonderful example of a team effort.

Look Ahead Plan: Aiming to make our last few EVAs fantastic ones and get in some great stargazing before we leave.

Anomalies in Work: None

Weather: Freezing in the morning, a bit warmer throughout midday and the afternoon. Calm with little wind and partly cloudy.

Crew Physical Status: All crew are physically healthy and safe.

EVA: See EVA Report.

Reports to be filed: Sol Summary Report, EVA Report, EVA Request, Photo Report, Commander’s Report, GreenHab Report, Astronomy Report

Support Requested: Please add Doug Campbell, incoming with Crew 193, to the Mission Support list

Maintenance request: See Operations Report

Astronomy Report – April 18th

Astronomy Report

Name: Victoria Varone

Crew: 192

Date: 18 April 2018

Sky Conditions: Partly Cloudy

Wind Conditions: Calm

Observation Start Time: 3:30 PM

Observation End Time: 4:00 PM

Summary: A quick scan and survey of the Sun’s surface provided a great view and a short telescope lesson for two of my crewmates.

Objects Viewed: Sun

Problems Encountered: A few clouds.

Greenhab Report – April 18th

GreenHab Report
Name: Andreea Radulescu
Date: 18 Apr 2018

Environmental control:

Ambient with door open few hours
Shade cloth: on
Fan: not on.

Average temperatures: 20.1℃
Low 13.6℃
High 53.2℃

Hours of supplemental light: None

Changes to crops: None

Daily water usage for crops: 9 gal

Time(s) of watering for crops: Twice: In the morning around 8:30 am and afternoon 6:30pm.

Moringa research observations: None

Changes to research plants: No

Aquaponics: None

Narrative: No significant changes. Picked another few baby carrots.

Support/supplies needed: None

EVA Report – April 18th

EVA Report

EVA #14

18 Apr 2018 – Sol 10

Crew members: Richard Blakeman, Ashok Narayanamoorthi, Victoria Varone
Objectives:
– Exploration reconnaissance of Lith Canyon Region for geology and terrain accessibility.
– Continue research on anti-fogging treatments by switching from two-piece spacesuits to one-piece spacesuits.
– Test hydration pack

Operations:
0800-1200

Narrative:

Performed an exploratory reconnaissance of Lith Canyon Region (Route 1106) to conduct geologic survey and collect imagery. Additionally, evaluated spacesuit visors for fogging study as well as conducted operational field testing of an experimental hydration system. Conducted preliminary terrain reconnaissance of Murphy Canyon region to pass vital information to upcoming EVA for follow up crew.

Vehicles
Rovers Spirit, Deimos, and Curiosity

__________

EVA #15

18 April 2018 – Sol 10

Crew Members: Joe Dituri, Andreea Radulescu

Objectives: Take Lowell Rd till we got to the base of 1106. We the proceeded south and walked the ridge line until the end.

Operations:
1400-1800

Narrative: This is absolutely the most beautiful part of the exploration. We walked to the end of the ridge…made our way down to the basin…Then proceeded down the water spill way which was another 30-50 feet down. Proceeding south we followed the water spillway until it turned almost due west…then proceeded north and then proceeded NW up the water spillway. We basically encircled Murphy’s Canyon as marked on the map in total. Almost 5 miles of hiking and several hundred foot change in elevation.

Vehicles
Rover Opportunity with two crew in it